1.find the vector form of the equation of the plane that passes through the point po=(1,-2,3) and has normal vector n=<3,1,-1>.

2.find an equation for the plane that contains the line x=-1+3t,y=5+3t,z=2+t and is parallel to the line of intersection of the planes x-2(y-1)+3z=-1 and y=-2x-1=0.


1
Expert's answer
2021-07-01T11:40:12-0400

1. P(1, -2, 3)


normal vector n = 3i + j - k



Position vector of point P = i - 2j + 3k


The vector form of a plane that consists of a vector r0 and a normal vector n is given by



n . ( r - r0) = 0


So, the required equation of plane is



(3i + j - k ) . [r - ( i - 2j + 3k ) ] = 0









2.line x=-1+3t,y=5+3t,z=2+t


The direction ratios of the line are d1= <3, 3, 1> and the point on the line is r0 = <-1, 5, 2>


The point that lies on the line would also lie on the plane. So, the position vector of the point is

- i + 5j + 2k.


The plane is parallel to the line of intersection of the planes

 x - 2y + 3z = -3 and 2x + y = 1


To find the line of intersection of the given planes we let z = t.

So,

x - 2y = -3 -3t and 2x + y = 1


On solving the above equations we get the line as :


<x, y, z> = <"\dfrac{6}{5}", "\dfrac{-7}{5}", 0 > + t <"\dfrac{3}{5}, \dfrac{-6}{5}, 1" >



The direction ratios of the line of intersection of the planes are d2 = <"\dfrac{3}{5}, \dfrac{-6}{5}, 1" >



Let the normal vector of the plane be n.


So,


n = d1 X d2

n = "\begin{vmatrix}
 i & j & k\\
 3 & 3 &1\\
3/5&-6/5&1
\end{vmatrix}" = "\dfrac{21}{5} i - \dfrac{12}{5}j - \dfrac{27}{5}k"




Hence the equation of the plane is



n.(r - r0) = 0


( "\dfrac{21}{5} i - \dfrac{12}{5}j - \dfrac{27}{5}k" ) . [ r - (-i + 5j + 2k)] = 0